# DMVC-Tracker: Distributed Multi-Agent Trajectory Planning for Target Tracking Using Dynamic Buffered Voronoi and Inter-Visibility Cells

### v1

@@ Abstract
Aerial tracking, path planning for multiple mobile robots, distributed robot systems.
@@ I Introduction
Aerial target tracking has been widely applied in fields such as cinematography [1] and surveillance [2]. Although a single micro aerial vehicle (MAV) is usually employed in these applications, the utilization of multiple MAVs can bring benefits. For example, multiple views of actors captured by a team of MAVs provide movie directors with more footage [3, 4, 5]. Also, multi-agent tracking can be deployed in moving motion-capture systems [6, 7, 8], and a large number of cameras increases the accuracy of pose estimation.
@@
Figure 1: Multi-agent target tracking missions in static (top) and dynamic (bottom) environments.
In this letter, we present an online distributed trajectory planning algorithm that can generate a target-visible and safe trajectory. The key ideas of the proposed method are Dynamic Buffered Voronoi Cell (DBVC) and Dynamic Inter-Visibility Cell (DIVC). The DBVC is a time-varying, inter-collision-free region developed from Buffered Voronoi Cell (BVC) [9]. This time-varying region helps maintain consistent distances between the target and the agents, whereas the BVC can cause an agent to become stuck in a static space, making it difficult for the agent to follow the target. We design the DIVC to prevent inter-agent occlusion. Specifically, the DIVC ensures that the Line-of-Sight connecting the target and an agent does not collide with other agents. Based on the DBVC and the DIVC, the proposed planner generates a target-visible and safe trajectory using Bernstein polynomial motion primitives, in an improved way over our previous work [10]. By reducing the conservativeness in [10], our method achieves a higher success rate. The main contributions of this paper are as follows.
@@
  • A construction method of Dynamic Buffered Voronoi Cell (DBVC) and Dynamic Inter-Visibility Cell (DIVC) that impose inter-agent collision and inter-agent occlusion avoidance constraints, while adapting to the future movements of the target.
@@ II-C Target Tracking in Crowded Environments
Most trajectory planning research aims to ensure real-time performance. The works [14], [18], and [19] formulate non-convex optimization methods to generate a target-visible trajectory; however, as the number of obstacles increases, computation time quickly increases. The work [16] adopts the QP-based approach for fast calculation, but it cannot handle unstructured obstacles. To address this, this work builds upon our previous work [10], using a Bernstein polynomial motion-primitive to efficiently find a tracking trajectory among both unstructured-but-static and dynamic-but-structured obstacles. Furthermore, this work reduces the conservativeness of the method in [10], further improving the tracking success rate in challenging conditions, such as environments with numerous dynamic obstacles.
@@
$\mathcal{X}$ Configuration space.
$\mathcal{T}(t)$ Occupied space by the target.
@@ III Problem Formulation
Extending to a 3D setup is straightforward, but fixed flying altitude enables the acquisition of consistent images of the target. Therefore, we address the 2D problem in this paper. Since the drone must find feasible movements only on the $x-y$ plane, the mission becomes more challenging. Throughout this paper, the notations in Table I are used.
@@ III-B Trajectory Representation
Due to the virtue of differential flatness of the dynamics of various MAV platforms, such as quadrotors, we represent trajectories of all agents as polynomial functions of time $t$. Employing the Bernstein basis [20], we represent the trajectory of the $i$-th agent, $\textbf{x}^{i}_{c}(t)\in\mathbb{R}^{2},i\in\mathcal{I}_{c}$ as follows:
@@ III-D1 Collision Avoidance
For safety, all agents should not collide with obstacles and the target.
$\mathcal{B}(\textbf{x}_{c}^{i}(t),r_{c})\subset\mathcal{F}_{q}(t),\ \forall i\in\mathcal{I}_{c},\ \forall{t}\in[0,T]$ (3)
@@ III-D2 Occlusion Avoidance
To avoid target occlusion by obstacles, the Lines-of-Sight between the agents and the target should not intersect with any obstacles.
$\mathcal{L}(\textbf{x}_{c}^{i}(t),\textbf{x}_{q}(t))\subset\mathcal{F}_{d}(t),\ \forall i\in\mathcal{I}_{c},\ \forall{t}\in[0,T]$ (4)
@@ III-D4 Dynamical Limits
$\|\dot{\textbf{x}}_{c}^{i}(t)\|\leq v_{\max},\ \forall i\in\mathcal{I}_{c},\ \forall t\in[0,T]$ (6a)
$\|\ddot{\textbf{x}}_{c}^{i}(t)\|\leq a_{\max},\ \forall i\in\mathcal{I}_{c},\ \forall t\in[0,T]$ (6b)
$|\dot{\psi}_{c}^{i}(t)|\leq\dot{\psi}_{\max},\ \forall i\in\mathcal{I}_{c},\ \forall t\in[0,T].$ (6c)
III-D5 Inter-Agent Collision Avoidance
For safety, the tracker team should also avoid inter-agent collisions.
$\mathcal{B}(\textbf{x}_{c}^{i}(t),r_{c})\subset\mathcal{F}_{c}^{i}(t),\ \forall i\in\mathcal{I}_{c},\ \forall{t}\in[0,T]$ (7)
III-D6 Inter-Agent Occlusion Avoidance
To avoid inter-agent occlusions, we ensure that the Lines-of-Sight between each agent and the target do not intersect with other agents.
$\mathcal{L}(\textbf{x}_{c}^{i}(t),\textbf{x}_{q}(t))\subset\mathcal{F}_{c}^{i}(t),\ \forall i\in\mathcal{I}_{c},\ \forall{t}\in[0,T]$ (8)
@@ IV-A Dynamic Buffered Voronoi Cell
Buffered Voronoi Cell (BVC) [9] has been used in various research as a technique for reciprocal collision avoidance in multi-agent trajectory planning. Since the BVC is static during the planning horizon, agents can be confined in the static cell while the target moves, leading to situations where the agents miss the target. Accordingly, we consider the inter-collision-free region that varies with the target’s future trajectory. We define a Dynamic Buffered Voronoi Cell (DBVC) to avoid collisions between the $i$-th and $j$-th agents as follows:
@@
$\frac{\textbf{x}_{c0}^{i}+\textbf{x}_{c0}^{j}}{2}\bigg{)}+r_{c}\|\textbf{x}_{c0}^{j}-\textbf{x}_{c0}^{i}\|\leq 0\bigg{\}},$ (9a)
@@
$\frac{\textbf{x}_{c0}^{j}+\textbf{x}_{c0}^{i}}{2}\bigg{)}+r_{c}\|\textbf{x}_{c0}^{i}-\textbf{x}_{c0}^{j}\|\leq 0\bigg{\}}$ (9b)
@@ IV-B Dynamic Inter-Visibility Cell
Similar to the Dynamic Buffered Voronoi Cell (DBVC), we define a Dynamic Inter-Visibility Cell (DIVC) that moves in accordance with the target’s future trajectory and prevents inter-agent occlusion between the $i$-th and $j$-th agents.
We design the DIVC by dividing the cases where an angle formed by the two Lines-of-Sight, $\mathcal{L}(\textbf{x}_{c0}^{i},\textbf{x}_{q0})$ and $\mathcal{L}(\textbf{x}_{c0}^{j},\textbf{x}_{q0})$, is either obtuse or acute. Also, depending on whether the relative configuration of current positions of the target, the $i$-th and $j$-th agents, i.e. $\textbf{x}_{q0}$, $\textbf{x}_{c0}^{i}$, and $\textbf{x}_{c0}^{j}$.
$z^{ij}=\text{sign}\bigg{(}\det\bigg{(}\begin{bmatrix}(\textbf{x}_{c0}^{j}-\textbf{x}_{q0})^{\top}\\ (\textbf{x}_{c0}^{i}-\textbf{x}_{q0})^{\top}\end{bmatrix}\bigg{)}\bigg{)}$ (12)
$z^{ij}$ is a variable determined by the sign function. Fig. 2 illustrates the cases where $z^{ij}=-1$.
@@ IV-B1 Obtuse case
Let $\textit{V}_{i}$, $\textit{V}_{j}$, and $\textit{V}_{Q}$ be points at $\textbf{x}_{c0}^{i}$, $\textbf{x}_{c0}^{j}$, and $\textbf{x}_{q0}$, respectively. Then, let points $V_{oi}$ and $V_{oj}$ be the points on line segments $\overline{V_{i}V_{Q}}$ and $\overline{V_{j}V_{Q}}$ that are each at a distance $\alpha_{o}^{ij}r_{c}$ ($\alpha_{o}^{ij}\geq 1$) from $V_{Q}$. We draw lines (red in Fig. 2(a)) that are perpendicular to $\overline{V_{i}V_{Q}}$ and $\overline{V_{j}V_{Q}}$, crossing $V_{oi}$ and $V_{oj}$, and we call half-spaces made by the lines $\mathcal{H}_{o1}^{ij}$ and $\mathcal{H}_{o1}^{ji}$, respectively. In addition, we draw the rays (blue in Fig. 2(a)) starting from $V_{oi}$ and $V_{oj}$, which are parallel to the lines that pass through $V_{Q}$ and are tangential to the balls, $\mathcal{B}(\textit{V}_{oj},r_{c})$ and $\mathcal{B}(\textit{V}_{oi},r_{c})$. Half-spaces made by the rays are represented as $\mathcal{H}_{o2}^{ij}$ and $\mathcal{H}_{o2}^{ji}$.
To satisfy the conditions that $V_{oi}$ and $V_{oj}$ are on $\overline{V_{i}V_{Q}}$ and $\overline{V_{j}V_{Q}}$, and the rays do not intersect, $\alpha_{o}^{ij}$ should be in the following range.
$&1\leq\alpha_{o}^{ij}\leq\min\Bigg{(}\frac{1}{r_{c}}\min(\|\textbf{x}_{c0}^{i}-\textbf{x}_{q0}\|,\|\textbf{x}_{c0}^{j}-\textbf{x}_{q0}\|),\\ &\quad\quad\quad\quad\quad\quad\sqrt{\frac{2}{1-\frac{(\textbf{x}_{c0}^{i}-\textbf{x}_{q0})^{\top}(\textbf{x}_{c0}^{j}-\textbf{x}_{q0})}{\|\textbf{x}_{c0}^{i}-\textbf{x}_{q0}\|\|\textbf{x}_{c0}^{j}-\textbf{x}_{q0}\|}}}\ \Bigg{)}\end{aligned}$ (13)
@@
The above inequalities mean that the distances between neighboring agents and Line-of-Sight connecting the tracker and the target are always greater than the size of the trackers $r_{c}$; therefore, the inter-occlusion constraints (8) are satisfied. To make the inter-occlusion-free regions move with the target, we maintain the shapes of $\mathcal{H}_{o1}^{ij}\cap\mathcal{H}_{o2}^{ij}$ and $\mathcal{H}_{o1}^{ji}\cap\mathcal{H}_{o2}^{ji}$ and translate them by $\textbf{x}_{q}(t)-\textbf{x}_{q0}$, equivalent to the amount of target’s movement. The moving half-spaces are represented as $\mathcal{H}_{o1}^{ij}(t)$, $\mathcal{H}_{o2}^{ij}(t)$, $\mathcal{H}_{o1}^{ji}(t)$, and $\mathcal{H}_{o2}^{ji}(t)$. Their mathematical expressions are simplified as follows, and $\mathcal{H}_{o1}^{ji}(t)$ and $\mathcal{H}_{o2}^{ji}(t)$ can be constructed by exchanging the indices $i$ and $j$ in (15).
$\mathcal{H}_{o1}^{ij}(t)=$$\{\textbf{x}(t)\in\mathbb{R}^{2}|(\textbf{x}_{c0}^{i}-\textbf{x}_{q0})^{\top}(\textbf{x}(t)-\textbf{x}_{q}(t))-$
$\ \alpha_{o}^{ij}r_{c}\|\textbf{x}_{c0}^{i}-\textbf{x}_{q0}\|\geq 0\},$ (15a)
$\mathcal{H}_{o2}^{ij}(t)=$$\bigg{\{}\textbf{x}(t)\in\mathbb{R}^{2}|\begin{bmatrix}z^{ij}\sin({\theta}_{cq}^{j}+z^{ij}\theta_{o}^{ij})\\ -z^{ij}\cos({\theta}_{cq}^{j}+z^{ij}\theta_{o}^{ij})\end{bmatrix}^{\top}\bigg{(}\textbf{x}(t)-$
$\ \textbf{x}_{q}(t)-\alpha_{o}^{ij}r_{c}\frac{(\textbf{x}_{c0}^{i}-\textbf{x}_{q0})}{\|\textbf{x}_{c0}^{i}-\textbf{x}_{q0}\|}\bigg{)}\leq 0\bigg{\}}$ (15b)
where $\alpha_{o}^{ij}$, $\theta_{cq}^{i}$, $\theta_{cq}^{j}$, and $\theta_{o}^{ij}$ are defined as follows.
$\begin{bmatrix}\cos(\theta_{cq}^{i})\\ \sin(\theta_{cq}^{i})\end{bmatrix}=\frac{(\textbf{x}_{c0}^{i}-\textbf{x}_{q0})}{\|\textbf{x}_{c0}^{i}-\textbf{x}_{q0}\|},\ \begin{bmatrix}\cos(\theta_{cq}^{j})\\ \sin(\theta_{cq}^{j})\end{bmatrix}=\frac{(\textbf{x}_{c0}^{j}-\textbf{x}_{q0})}{\|\textbf{x}_{c0}^{j}-\textbf{x}_{q0}\|},$ (16a)
$\sin(\theta_{o}^{ij})=\frac{1}{\alpha_{o}^{ij}},\ \cos(\theta_{o}^{ij})=\sqrt{1-\sin^{2}(\theta_{o}^{ij})}$ (16b)

### v2

@@ Abstract
Path planning for multiple mobile robots, distributed robot systems, aerial tracking.
@@ I Introduction
Aerial target tracking has been widely applied in fields such as cinematography and surveillance. Although a single micro aerial vehicle (MAV) is usually employed in these applications, the utilization of multiple MAVs can bring benefits. For example, multiple views of actors captured by a team of MAVs provide movie directors with more footage [1, 2, 3]. Also, multi-agent tracking can be deployed in moving motion-capture systems [4, 5, 6], and a large number of cameras increases the accuracy of pose estimation.
@@
Figure 1: Hardware demonstration of multi-agent target tracking.
In this letter, we present an online distributed trajectory planning algorithm that can generate a target-visible and safe trajectory. The key ideas of the proposed method are Dynamic Buffered Voronoi Cell (DBVC) and Dynamic Inter-Visibility Cell (DIVC). The DBVC is a time-varying, inter-collision-free region developed from Buffered Voronoi Cell (BVC) [7]. This time-varying region helps maintain consistent distances between the target and the agents, whereas the BVC can cause an agent to become stuck in a static space, making it difficult for the agent to follow the target. We design the DIVC to prevent inter-agent occlusion. Specifically, the DIVC ensures that the Line-of-Sight connecting the target and an agent does not collide with other agents. Based on the DBVC and the DIVC, the proposed planner generates a target-visible and safe trajectory using Bernstein polynomial motion primitives. We employ a sample-check-select strategy and leverage properties of Bernstein polynomials to speed up the calculation, enabling us effectively to tackle complex tracking problems. This paper reduces the conservativeness of the check method from our previous work [8] to achieve a higher success rate. The main contributions of this paper are as follows.
@@
  • Dynamic Buffered Voronoi Cell (DBVC) and Dynamic Inter-Visibility Cell (DIVC) that impose inter-agent collision and inter-agent occlusion avoidance constraints, while adapting to the future movements of the target.
@@ II-C Target Tracking in Crowded Environments
Most trajectory planning research aims to ensure real-time performance. The works [12], [16], and [17] formulate non-convex optimization methods to generate a target-visible trajectory; however, as the number of obstacles increases, computation time quickly increases. To address this, this work employs Bernstein polynomial motion primitives to efficiently find a tracking trajectory by leveraging the properties of Bernstein polynomials. This work reduces the conservativeness of the approach in our previous work [8], further improving the tracking success rate in challenging conditions, such as environments with numerous dynamic obstacles.
@@
$\mathcal{X}$, $\mathcal{T}(t)$ Environment, Occupied space by the target.
@@ III Problem Formulation
For consistent image acquisition of the target, we fix the flying altitude, which means we address the 2D problem. Unlike methods that allow drones to move obstacle-free high altitudes [18, 19], since the drone must find feasible movements only on the $x$-$y$ plane, our problem is more challenging. Throughout this paper, the notations in Table I are used.
@@ III-B Trajectory Representation
Due to the virtue of differential flatness of the dynamics of various MAV platforms, such as quadrotors, we represent trajectories of all agents as polynomial functions of time $t$. Employing the Bernstein basis, we represent the trajectory of the $i$-th agent, $\textbf{x}^{i}_{c}(t)\in\mathbb{R}^{2},i\in\mathcal{I}_{c}$ as follows:
@@ III-D1 Collision Avoidance
For safety, all agents should not collide with obstacles, the target, and the other agents.
$\mathcal{B}(\textbf{x}_{c}^{i}(t),r_{c})\subset\mathcal{F}_{q}(t)\cap\mathcal{F}_{c}^{i}(t),\ \forall i\in\mathcal{I}_{c},\ \forall{t}\in[0,T]$ (3)
@@ III-D2 Occlusion Avoidance
To avoid target occlusion caused by obstacles or neighboring agents, the Lines-of-Sight between the agents and the target should not intersect with any obstacles and other agents.
$\mathcal{L}(\textbf{x}_{c}^{i}(t),\textbf{x}_{q}(t))\subset\mathcal{F}_{d}(t)\cap\mathcal{F}_{c}^{i}(t),\ \forall i\in\mathcal{I}_{c},\ \forall{t}\in[0,T]$ (4)
@@ III-D4 Dynamical Limits
$\|\dot{\textbf{x}}_{c}^{i}(t)\|\leq v_{\max},\ \|\ddot{\textbf{x}}_{c}^{i}(t)\|\leq a_{\max},$ $|\dot{\psi}_{c}^{i}(t)|\leq\dot{\psi}_{\max},$ (6)
$\forall i\in\mathcal{I}_{c},\ \forall t\in[0,T]$
@@ IV-A Dynamic Buffered Voronoi Cell
We consider the inter-collision-free region that moves along the target’s predicted trajectory. We define a Dynamic Buffered Voronoi Cell (DBVC) to avoid collisions between the $i$-th and $j$-th agents as follows:
@@
$\frac{\textbf{x}_{c0}^{i}+\textbf{x}_{c0}^{j}}{2}\Big{)}+r_{c}\|\textbf{x}_{c0}^{j}-\textbf{x}_{c0}^{i}\|\leq 0\Big{\}},$ (7a)
@@
$\frac{\textbf{x}_{c0}^{j}+\textbf{x}_{c0}^{i}}{2}\Big{)}+r_{c}\|\textbf{x}_{c0}^{i}-\textbf{x}_{c0}^{j}\|\leq 0\Big{\}}$ (7b)
@@ IV-B Dynamic Inter-Visibility Cell
Similar to the Dynamic Buffered Voronoi Cell (DBVC), we define a Dynamic Inter-Visibility Cell (DIVC) that moves in accordance with the target’s future trajectory and prevents inter-agent occlusion between the $i$-th and $j$-th agents. We design the DIVC by dividing the cases where an angle formed by the two Lines-of-Sight, $\mathcal{L}(\textbf{x}_{c0}^{i},\textbf{x}_{q0})$ and $\mathcal{L}(\textbf{x}_{c0}^{j},\textbf{x}_{q0})$, is either obtuse or acute.
@@ IV-B1 Obtuse case
Let $\textit{V}_{i}$, $\textit{V}_{j}$, and $\textit{V}_{Q}$ be points at $\textbf{x}_{c0}^{i}$, $\textbf{x}_{c0}^{j}$, and $\textbf{x}_{q0}$, respectively. Then, let points $V_{oi}$ and $V_{oj}$ be the points on line segments $\overline{V_{i}V_{Q}}$ and $\overline{V_{j}V_{Q}}$ that are each at a distance $\alpha_{o}^{ij}r_{c}$ ($\alpha_{o}^{ij}\geq 1$) from $V_{Q}$. We draw lines (red in Fig. 2(a)) that are perpendicular to $\overline{V_{i}V_{Q}}$ and $\overline{V_{j}V_{Q}}$, crossing $V_{oi}$ and $V_{oj}$, and we call half-spaces made by the lines $\mathcal{H}_{o1}^{ij}$ and $\mathcal{H}_{o1}^{ji}$, respectively. In addition, we draw the rays (blue in Fig. 2(a)) starting from $V_{oi}$ and $V_{oj}$, which are parallel to the lines that pass through $V_{Q}$ and are tangential to the balls, $\mathcal{B}(\textit{V}_{oj},r_{c})$ and $\mathcal{B}(\textit{V}_{oi},r_{c})$. Half-spaces made by the rays are represented as $\mathcal{H}_{o2}^{ij}$ and $\mathcal{H}_{o2}^{ji}$. To satisfy the conditions that $V_{oi}$ and $V_{oj}$ are on $\overline{V_{i}V_{Q}}$ and $\overline{V_{j}V_{Q}}$, and the rays do not intersect, $\alpha_{o}^{ij}$ should be in the following range.
$1\leq\alpha_{o}^{ij}\leq\min\Bigg{(}\frac{\min(\|\textbf{x}_{cq0}^{i}\|,\|\textbf{x}_{cq0}^{j}\|)}{r_{c}},\sqrt{\frac{2}{1-(\textbf{n}_{cq0}^{i})^{\top}\textbf{n}_{cq0}^{j}}}\Bigg{)}$ (10)
$\text{where}\ \textbf{x}_{cq0}^{i}=\textbf{x}_{c0}^{i}-\textbf{x}_{q0},\ \textbf{x}_{cq0}^{j}=\textbf{x}_{c0}^{j}-\textbf{x}_{q0},$
$\quad\quad\ \ \textbf{n}_{cq0}^{i}=\textbf{x}_{cq0}^{i}/\|\textbf{x}_{cq0}^{i}\|,\ \textbf{n}_{cq0}^{j}=\textbf{x}_{cq0}^{j}/\|\textbf{x}_{cq0}^{j}\|$
@@
The above inequalities mean that the distances between neighboring agents and Line-of-Sight connecting the tracker and the target are always greater than the size of the trackers $r_{c}$; therefore, the inter-occlusion constraints are satisfied. To make the inter-occlusion-free regions move with the target, we maintain the shapes of $\mathcal{H}_{o1}^{ij}\cap\mathcal{H}_{o2}^{ij}$ and $\mathcal{H}_{o1}^{ji}\cap\mathcal{H}_{o2}^{ji}$ and translate them by $\textbf{x}_{q}(t)-\textbf{x}_{q0}$, equivalent to the amount of target’s movement. The moving half-spaces are represented as $\mathcal{H}_{o1}^{ij}(t)$, $\mathcal{H}_{o2}^{ij}(t)$, $\mathcal{H}_{o1}^{ji}(t)$, and $\mathcal{H}_{o2}^{ji}(t)$. Their mathematical expressions are simplified as follows, and $\mathcal{H}_{o1}^{ji}(t)$ and $\mathcal{H}_{o2}^{ji}(t)$ can be constructed by exchanging the indices $i$ and $j$ in (12).
$\mathcal{H}_{o1}^{ij}(t)=$$\{\textbf{x}(t)\in\mathbb{R}^{2}|(\textbf{n}_{cq0}^{i})^{\top}(\textbf{x}(t)-\textbf{x}_{q}(t))-\alpha_{o}^{ij}r_{c}\geq 0\},$ (12a)
$\mathcal{H}_{o2}^{ij}(t)=$$\bigg{\{}\textbf{x}(t)\in\mathbb{R}^{2}|\begin{bmatrix}z^{ij}\sin({\theta}_{cq}^{j}+z^{ij}\theta_{o}^{ij})\\ -z^{ij}\cos({\theta}_{cq}^{j}+z^{ij}\theta_{o}^{ij})\end{bmatrix}^{\top}(\textbf{x}(t)-$
$\ \textbf{x}_{q}(t)-\alpha_{o}^{ij}r_{c}\textbf{n}_{cq0}^{i})\leq 0\bigg{\}}$ (12b)
where $z^{ij}$, $\theta_{cq}^{i}$, $\theta_{cq}^{j}$, and $\theta_{o}^{ij}$ are defined as follows.
@@
$z^{ij}=\text{sign}\big{(}\det\big{(}[(\textbf{x}_{cq0}^{j})^{\top};(\textbf{x}_{cq0}^{i})^{\top}]\big{)}\big{)},$ (13)
$\begin{bmatrix}\cos(\theta_{cq}^{i})\\ \sin(\theta_{cq}^{i})\end{bmatrix}=\textbf{n}_{cq0}^{i},\ \begin{bmatrix}\cos(\theta_{cq}^{j})\\ \sin(\theta_{cq}^{j})\end{bmatrix}=\textbf{n}_{cq0}^{j},$
$\sin(\theta_{o}^{ij})=(\alpha_{o}^{ij})^{-1},\ \cos(\theta_{o}^{ij})=\sqrt{1-\sin^{2}(\theta_{o}^{ij})}$